Use the linear equation of state of seawater to calculate the density of freshwater (0 salinity) and ocean water (35 salinity) at 15C and zero pressure (relative to sea level pressure).

where p 1027 kg/m3, T-10 °C, S-35 psu, a--0.15 kg/m3/°C, b-0.78 kg/m3/psu, k-4.5x10-3 kg/m3/dbar
Answer #1

Given equation : p - pr = a * ( T - Tr ) + b* ( S - Sr ) + k * P

Density of freshwater p = a * ( T - Tr )+b* ( S - Sr ) + k * P + pr

Density of freshwater p = - 0.15 * ( 15 - 10 ) + 0.78 ( 0 - 35 ) + 0.0045 * 0 + 1027

Density of freshwater p = -0.75 - 27 .3 + 0 + 1027

Density of freshwater p = 998.95 kg / m3

Density of ocean water p = a * ( T - Tr )+b* ( S - Sr ) + k * P + pr

Density of ocean water p = - 0.15 * ( 15 - 10 ) + 0.78 ( 35 - 35 ) + 0.0045 * 0 + 1027

Density of ocean water p = -0.75 + 1027 = 1026.25 kg / m3